OpenStudy (anonymous):

A person weighs 500 N. If he falls a vertical distance of 15 meters into a pile of hay, what is the average force exerted by the hay if the person stops in 0.12 seconds?

OpenStudy (anonymous):

take g=10, we get m=50 use the use kinematic for falling body to calculate the velocity \[ v = \sqrt{2gh} \] and \[ v = a t \] so we get \[ F = ma = \frac{m}{t}\sqrt{2gh} \]